OpenStudy (anonymous):

subtract: (9+2i)-(8-5i)

OpenStudy (saifoo.khan):

1 + 7 i

OpenStudy (anonymous):

1-3i

OpenStudy (anonymous):

9-8+2i-(-5i)=1+2i+5i = 1+ 7i